# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ III-D Efficient Linear Fast Attention Mechanism
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &=\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\textit{{V}}\\ &={\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\left[\bar{\textit{{V}}},\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right],\\ \end{split}$
@@
where $\odot$ denotes the Hadamard product [36], ${\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\tilde{\textit{{V}}}$ is equivalent to $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)$, and $\textit{{Z}}\in\mathbb{R}^{m\times m}$ is a matrix as follows,
@@
with $\otimes$ denoting the Kronecker product [36] and $\textit{{U}}_{(m/b)}\in\mathbb{R}^{(m/b)\times(m/b)}$ being the all-ones matrix. It is obvious to see that
@@
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi},\end{split}$ (9)
where $\xi=2$ denotes the spectral norm and $\xi=F$ denotes the Frobenius norm. For given two matrices $\textit{{A}}\in\mathbb{R}^{m\times n}$ and $\textit{{B}}\in\mathbb{R}^{n\times p}$, it follows from [37, Section 2.3.1] that $\|\textit{{A}}\textit{{B}}\|_{\xi}\leq\|\textit{{A}}\|_{\xi}\|\textit{{B}}\|_{\xi}$.
@@
Let $\emph{{{U}}}_{m}\in\mathbb{R}^{m\times m}$ be an all-ones matrix. For any two vectors $\textit{{q}}\in\mathbb{R}^{c}$ from $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{k}}\in\mathbb{R}^{c}$ from $\textit{{K}}\in\mathbb{R}^{m\times c}$, let $\mathfrak{M}>0$ and $\mathscr{M}>0$ be the maximum and minimum of ${\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right)\right)$, respectively, and $q_{i}$ and $k_{i}$ are the elements at position $i$ of vectors q and k, respectively. If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
@@
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
@@
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
@@
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Before proving Theorem 1, we introduce the following lemma (see [5]).
Lemma 1.
For any two vectors $\textit{{q}}\in\mathbb{R}^{c}$ from $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{k}}\in\mathbb{R}^{c}$ from $\textit{{K}}\in\mathbb{R}^{m\times c}$, let
$\begin{split}D_{\textit{{q}},\textit{{k}}}&=\max_{i=1,2,\dots,c}{\rm exp}(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right))>0,\\ d_{\textit{{q}},\textit{{k}}}&=\min_{i=1,2,\dots,c}{\rm exp}(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right))>0,\end{split}$
@@
where $q_{i}$ and $k_{i}$ are the elements at position $i$ of vectors q and k, respectively. The following inequalities hold,
$\begin{split}\frac{{\rm exp}\left(\textit{{q}}\right){\rm exp}\left(\textit{{k}}\right)^{\top}}{{\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}\right)}&\leq\frac{c}{d_{\textit{{q}},\textit{{k}}}{\rm exp}\left(-0.5\right)},\\ \frac{{\rm exp}\left(\textit{{q}}\right){\rm exp}\left(\textit{{k}}\right)^{\top}}{{\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}\right)}&\geq\frac{c}{D_{\textit{{q}},\textit{{k}}}{\rm exp}\left(-0.5\right)}.\end{split}$
The following corollary is easily obtained from Lemma 1.
Corollary 1.
For two matrices $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{K}}\in\mathbb{R}^{m\times c}$, let
@@
$\begin{split}\mathfrak{M}&=\max_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c}{\rm exp}(\textit{{q}}_{i_{1}}\textit{{k}}_{i_{2}}^{\top}+0.5-\left(q_{i_{1}j}+k_{i_{2}j}\right))>0,\\ \mathscr{M}&=\min_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c}{\rm exp}(\textit{{q}}_{i_{1}}\textit{{k}}_{i_{2}}^{\top}+0.5-\left(q_{i_{1}j}+k_{i_{2}j}\right))>0,\end{split}$
@@
where $q_{i_{1}j}$ and $k_{i_{2}j}$ are the elements at position $j$ of the vector $\textit{{q}}_{i_{1}}=\textit{{Q}}(i_{1},:)$ and the vector $\textit{{k}}_{i_{2}}=\textit{{K}}(i_{2},:)$, respectively. The following inequalities hold,
@@
$\begin{split}\frac{\left\|{\rm exp}\left(\textit{{Q}}\right){\rm exp}\left(\textit{{K}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\right\|_{\xi}}&\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)},\\ \frac{\left\|{\rm exp}\left(\textit{{Q}}\right){\rm exp}\left(\textit{{K}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\right\|_{\xi}}&\geq\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}.\end{split}$
@@
Following from Corollary 1, if $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
@@
We now consider the upper bound for the right-hand side of Inequality (9). If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,

### v4

@@ Appendix A Approximation Error Bound Analysis
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}=\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\textit{{V}}\\ &={\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\left[\bar{\textit{{V}}},\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\quad\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right],\\ \end{split}$
@@
where $\odot$ denotes the Hadamard product (Horn and Johnson, 2012), ${\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\tilde{\textit{{V}}}$ is equivalent to $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)$, and $\textit{{Z}}\in\mathbb{R}^{m\times m}$ is a matrix as follows,
@@
with $\otimes$ denoting the Kronecker product (Horn and Johnson, 2012) and $\textit{{U}}_{(m/b)}\in\mathbb{R}^{(m/b)\times(m/b)}$ being the all-ones matrix. It is obvious to see that
@@
(12) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi},\end{split}$
where $\xi=2$ denotes the spectral norm and $\xi=F$ denotes the Frobenius norm. For given two matrices $\textit{{A}}\in\mathbb{R}^{m\times n}$ and $\textit{{B}}\in\mathbb{R}^{n\times p}$, it follows from (Golub and Van Loan, 2013) [Section 2.3.1] that $\|\textit{{A}}\textit{{B}}\|_{\xi}\leq\|\textit{{A}}\|_{\xi}\|\textit{{B}}\|_{\xi}$.
@@
Let $\emph{{{U}}}_{m}\in\mathbb{R}^{m\times m}$ be an all-ones matrix. For any two vectors $\bar{\textit{{q}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{k}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let $\mathfrak{M}>0$ and $\mathscr{M}>0$ be the maximum and minimum of ${\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)$, respectively, and $\bar{q}_{i}$ and $\bar{k}_{i}$ are the elements at position $i$ of vectors $\bar{\textit{{q}}}$ and $\bar{\textit{{k}}}$, respectively. If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
@@
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
@@
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
@@
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Before proving Theorem A.1, we introduce the following lemma (see (Wu et al., 2024)).
Lemma A.2.
For any two vectors $\bar{\textit{{q}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{k}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let
$\begin{split}D_{\bar{\textit{{q}}},\bar{\textit{{k}}}}&=\max_{i=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)>0,\\ d_{\bar{\textit{{q}}},\bar{\textit{{k}}}}&=\min_{i=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)>0,\end{split}$
@@
where $\bar{q}_{i}$ and $\bar{k}_{i}$ are the elements at position $i$ of vectors $\bar{\textit{{q}}}$ and $\bar{\textit{{k}}}$, respectively. The following inequalities hold,
$\begin{split}\frac{{\rm exp}\left(\bar{\textit{{q}}}\right){\rm exp}\left(\bar{\textit{{k}}}\right)^{\top}}{{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}\right)}&\leq\frac{c_{1}}{d_{\bar{\textit{{q}}},\bar{\textit{{k}}}}{\rm exp}\left(-0.5\right)},\\ \frac{{\rm exp}\left(\bar{\textit{{q}}}\right){\rm exp}\left(\bar{\textit{{k}}}\right)^{\top}}{{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}\right)}&\geq\frac{c_{1}}{D_{\bar{\textit{{q}}},\bar{\textit{{k}}}}{\rm exp}\left(-0.5\right)}.\end{split}$
The following corollary is easily obtained from Lemma A.2.
Corollary A.3.
For two matrices $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let
@@
$\begin{split}\mathfrak{M}&=\max_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}_{i_{1}}\bar{\textit{{k}}}_{i_{2}}^{\top}+0.5-\left(\bar{q}_{i_{1}j}+\bar{k}_{i_{2}j}\right)\right)>0,\\ \mathscr{M}&=\min_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}_{i_{1}}\bar{\textit{{k}}}_{i_{2}}^{\top}+0.5-\left(\bar{q}_{i_{1}j}+\bar{k}_{i_{2}j}\right)\right)>0,\end{split}$
@@
where $\bar{q}_{i_{1}j}$ and $\bar{k}_{i_{2}j}$ are the elements at position $j$ of the vector $\bar{\textit{{q}}}_{i_{1}}=\bar{\textit{{Q}}}(i_{1},:)$ and the vector $\bar{\textit{{k}}}_{i_{2}}=\bar{\textit{{K}}}(i_{2},:)$, respectively. The following inequalities hold,
@@
$\begin{split}\frac{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}}&\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)},\\ \frac{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}}&\geq\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}.\end{split}$
@@
Following from Corollary A.3, if $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
@@
We now consider the upper bound for the right-hand side of Inequality (12). If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,